OpenStudy (anonymous):

What is the solution set for |–2x + 16| – 9 = –3

OpenStudy (anonymous):

use -(-2x+16) and (2x+16) and solve the eqn both ways

OpenStudy (anonymous):

l-2x + 16l = 6 => -2x + 16 = 6 , -2x + 16 = -6 => x = 5 , x = 11

OpenStudy (anonymous):

I agree with abhinareddy

OpenStudy (anonymous):

pls .....it is abhinayreddy

OpenStudy (anonymous):

abhina had a shorter way of doing it